A triangle has sides measured 5 inches , 12 inches , and 13 inches respectively ls this triangle a right triangle ? How do you know?
Answer by lwsshak3(11628)

A triangle has sides measured 5 inches , 12 inches , and 13 inches respectively ls this triangle a right triangle ? How do you know?
***
For a right triangle, sum of the squares of the two shorter sides=square of the longest side.
5^2+12^2=25+144=169=13^2
Comment: This is one of the right triangles you should immediately recognize when you see it.